OpenStudy (anonymous):

verify solution for : log(57x)= 2 + log (x-2)

OpenStudy (anonymous):

i get x= 200/43 i am not sure if that is correct

hartnn (hartnn):

how did you get that ?

OpenStudy (anonymous):

\[\log(57x)= 2 + \log (x-2) \] \[\log(57x)-\log(x-2)=\log (\frac{57x}{x-2})\] \[\log (\frac{57x}{x-2})=2\] \[10^2=\frac{57x}{x-2}\] \[100x-200=57x\] \[43x=200\] \[x=\frac{200}{43}\]

OpenStudy (anonymous):

exactly! thanks thats how I got it

OpenStudy (anonymous):

Good job mate.